# Analysis of a finite element method for second order uniformly elliptic PDEs in non-divergence form

### v1

@@ Abstract.
We propose one finite element method for both second order linear uniformly elliptic PDE in non-divergence form and the elliptic Hamilton-Jacobi-Bellman (HJB) equation. For the linear elliptic PDE in non-divergence form, we consider two scenarios of the matrix coefficient matrix $A$. One is $A$ is uniformly continuous. The other is $A$ is discontinuous but $\gamma A$ is dominated by $I_{d}$ where $\gamma$ is a positive weight function.
We prove that optimal convergence in discrete $W^{2,p}$-norm of the numerical approximation to the strong solution for $1<p\leq 2$ on convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$). If the domain is a two dimensional non-convex polygon, $p$ is valid in a neighbourhood of $\frac{4}{3}$. We also prove the well-posedness of strong solution in $W^{2,p}(\Omega)$ for both linear elliptic PDE in non-divergence form and the HJB equation for $1<p\leq 2$ on convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$) and for $p$ in an open interval starting from $1$ and including $\frac{4}{3}$ on two dimensional non-convex polygon. Furthermore, we relax the assumptions on the continuity of coefficients of the HJB equation, which have been widely used in literature.
@@ 1. Introduction
We mainly consider optimal convergence to the strong solution in $W_{h}^{2,p}(\Omega)$-norm (see (7)) of a finite element method for both second order linear uniformly elliptic PDE in non-divergence form (2) and the elliptic Hamilton-Jacobi-Bellman equation (3) on Lipschitz polyhedral domains in $\mathbb{R}^{d}$ ($d=2,3$). We consider the well-posedness of the strong solution in $W^{2,p}(\Omega)$ of (2) and (3) as well. The ranges of $p$ are described below:
@@
The linear elliptic PDE in non-divergence form is find the solution $u$ satisfying
We assume that $A=A(\boldsymbol{x})\in[L^{\infty}(\Omega)]^{d\times d}$ is a symmetric and uniformly positive definite matrix-valued function on $\Omega$, $\boldsymbol{b}\in[L^{\infty}(\Omega)]^{d}$, and $c\in L^{\infty}(\Omega)$ with $c\leq 0$.
The elliptic Hamilton-Jacobi-Bellman (HJB) equation is to find the solution $u$ satisfying
We further assume that there is a subset $E_{0}\subset\Omega$ with zero $d$-dimensional Lebesgue measure such that
@@
1.1. Literature review
There are several papers [1, 3, 6, 8, 12, 15, 19, 27, 29, 33, 32, 37] focusing on convergence to the viscosity solutions. In the Barles–Souganidis framework [1], monotonicity serves as a fundamental principle underpinning numerical schemes to ensure convergence to viscosity solutions of fully nonlinear partial differential equations. This concept is also applicable to the monotone finite difference methods for the HJB equation [3, 19]. A monotone finite element like scheme for the HJB equation was proposed in [6]. In [32], Nochetto and Zhang introduced a two-scale method based on the integro-differential formulation of (2), where a discrete Alexandroff–Bakelman–Pucci estimate was established. This two-scale method [32] was extended to the Isaacs equation [37]. However, up to our knowledge, all these numerical methods for convergence to the viscosity are low order. In addition, most of them have stencils which increase under mesh refinement, such that the corresponding linear system may not be sparse.
With respect to numerical methods for the convergence to the strong solutions, there are tree main categories.
  • (I)
    AA can be discontinuous but $\gamma A$ dominated by $I_{d}$ where $\gamma$ is a positive weight function. In [39], authors introduced and analyzed a Discontinuous Galerkin (DG) method for (2). The main assumptions are the coefficients $A$ is dominated by $I_{d}$ after being multiplied by a positive weight function $\gamma$ (Cordès condition) and the domain $\Omega$ is convex. The stability and optimal rate of convergence in discrete $H^{2}$-norm were proven in [39]. Following this direction, there are several papers [2, 21, 43]. The technique in [39] has been extended to the HJB equation (3) in [40, 22, 43] but with assumption that $A^{\alpha}\in C^{0}(\overline{\Omega}\times\Lambda)$ and the index set $\Lambda$ is a compact metric space. It seems that numerical methods along this direction may only be able to obtain optimal rate of convergence with respect to discrete $H^{2}$-type norm ($p=2$).
  • (II)
    AA is globally continuous. In [14], authors introduced and analyzed a $C^{0}$-conforming DG method for (2). The main assumptions are $A\in C^{0}(\overline{\Omega})$ and the domain $\Omega$ has $C^{1,1}$ boundary. They followed the approach in [24, Chapter $9$] to prove that their numerical method could achieve stability and optimal rate of convergence in discrete $W^{2,p}$-norm for any $1<p<+\infty$. However, this numerical method and its variant - a DG method in [17] can not deal with discontinuous coefficients $A$, since they have the jump of $A\nabla u_{h}$ along mesh interfaces ($u_{h}$ is the numerical solution). Therefore, the numerical methods in [14, 17] can not be applied for the HJB equation (3).
  • (III)
    Residual minimization methods. The main advantage of residual minimization methods [35, 41, 23] is that they can easily inherit the a priori estimate of PDE operators of (2) and (3). However, the minimal requirements of the regularity of true solution $u$ for these methods are highly affected by the choice of the norm to measure the residual in their designs. It is not clear which norm to choose for the two scenarios mentioned above.
A natural question is whether it is possible to design one numerical method for both linear elliptic PDE in non-divergence form and the elliptic HJB equation on Lipschitz polyhedra in $\mathbb{R}^{d}$ ($d=2,3$) and achieve optimal convergence in discrete $W^{2,p}$-norm to the strong solution with a wide range of $p$.
1.2. Main contributions
1.2.1. Contributions for (2)
@@
Theorem 3.4 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ and optimal convergence with respect to $W_{h}^{2,p}$-norm (see (7)) of the numerical solution of the FEM (14), in the scenario where the coefficient matrix $A\in C^{0}(\overline{\Omega})$. The results of well-posedness of (2) provided by [24, Theorem $9.15$] and optimal convergence of the FEM in [14] require $\partial\Omega$ is $C^{1,1}$. Actually, in neighbourhood of any point on $\partial\Omega$, the analysis in [24, Theorem $9.15$] eventually relies on [24, Lemma $9.12$]. The proof of [24, Lemma $9.12$] needs to “flatten” $\partial\Omega$ such that $\partial\Omega$ needs to be $C^{1,1}$. In contrast, Theorem 3.4 is valid on Lipschitz polyhedral domains (see (1) for the detailed description). Furthermore, the existence of strong solution is mainly proven by the discrete compactness - Lemma 3.8 and the stability of the FEM - Theorem 3.3, while the convergence analysis [14, Theorem 3.2] is based on [24, Theorem $9.15$]. Therefore, even for (2) with uniformly continuous coefficient matrix $A$, the numerical analysis in this paper is quite different from that in [14]. We refer to the statement at the beginning of Section 3 for detailed description of the methodologies used in this paper for (2) with uniformly continuous $A$.
@@
Theorem 3.5 and Theorem 3.6 provides well-posedness of strong solution in $W^{2,p}(\Omega)$ of (2) and optimal convergence with respect to $W_{h}^{2,p}$-norm of the numerical solution of the FEM (14), in the scenario where the coefficient matrix $A$ is discontinuous but $\gamma A$ is dominated by $I_{d}$, where the positive weight function $\gamma$ is introduced in (12). Since Theorem 3.5 and Theorem 3.6 are special examples of Theorem 4.1 and Theorem 4.3 respectively, the comparison below between Theorem 3.5, Theorem 4.3 and results in [40] for the HJB equation (3) is directly relevant to Theorem 3.5 and Theorem 3.6 as well.
1.2.2. Contributions for (3)
@@
For the elliptic HJB equation (3), the main theoretical results are Theorem 4.1 and Theorem 4.3. With (4,5) and the assumption (6), Theorem 4.1 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ of the elliptic HJB equation (3), while Theorem 4.3 provides optimal convergence of the FEM (16). Please see Remark 4.2 and Remark 4.4 for the special case $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$ and $c^{\alpha}=0$ for any $\alpha\in\Lambda$.
When the domain is convex and $p=2$, the condition (89) used in Theorem 4.3 is more restrictive than the Cordès condition used in [40]. It is due to two reasons. The first reason is that the constant of the discrete Miranda–Talenti inequality of the discrete Lapalcian operator $L_{I,h}$ associated with the FEM (16) is not known explicitly and probably bigger than $1$. The second reason is that we don’t want to include the parameter $\lambda$, which may not be known beforehand, in the design of the weight function $\gamma^{\alpha}$ (see the difference between (15) and the weight function used in [40]). Though it is more restrictive, Remark 4.5 shows that the condition (89) allows non-trivial examples. On the other hand, Theorem 4.3 is valid for a wide range of $p$ (see (1b)) and is valid on two dimensional non-convex polygon for a neighbourhood of $\frac{4}{3}$.
@@
We would like to emphasize that [40] assumes $A^{\alpha}\in C^{0}(\overline{\Omega}\times\Lambda)$ and the index set $\Lambda$ is a compact metric space, while [23, Proposition $4.3$] assumes that there are $\beta\in(0,1]$ and $0<C<+\infty$ such that $\|A^{\alpha}\|_{C^{0,\beta}(\overline{\Omega})}<C$ for any $\alpha\in\Lambda$ ([23, Proposition $4.3$] assumes $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$, $c^{\alpha}=0$, and $f^{\alpha}=f$). So, if the index set $\Lambda$ has a single element, the elliptic HJB equation (3) will not become the linear elliptic PDE in non-divergence form (2) with discontinuous matrix coefficient $A$. In contrast, the assumption (6) allows discontinuity of $A^{\alpha}$, $\boldsymbol{b}^{\alpha}$ and $c^{\alpha}$. By Lemma 4.3, we can conclude that the assumptions used by [40] and ([23, Proposition $4.3$] imply that the assumption (6) holds.
@@ 3. Linear elliptic PDE in non-divergence form
Theorem 3.4 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ and optimal convergence with respect to $W_{h}^{2,p}$-norm (see (7)) of the numerical solution of the FEM (14), in the scenario where the coefficient matrix $A\in C^{0}(\overline{\Omega})$. We explain the methodology to prove Theorem 3.4 by the following steps.
  • Step $1$. We prove Theorem 3.2 which give the following global $W^{2,p}$ estimate (22) of the linear elliptic PDE in non-divergence form (2). The most difficult part of the proof of Theorem 3.2 is to obtain (22) for $1<p\leq\frac{3}{2}$ if $\Omega$ is a convex polyhedra in $\mathbb{R}^{3}$, since the Alexandroff–Bakelman–Pucci estimate can not be applied immediately due to the fact $W^{2,p}(\Omega)$ can not be embedded into $C^{0}(\overline{\Omega})$. To overcome this difficulty, we provide Lemma 3.1 which shows the uniqueness of the solution of (2) in $W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$.
  • Step $2$. We provide Lemma 3.7 which give the inequality (38). The proof of Lemma 3.7 mimics that of [14, Lemma $3.4$]. However, due to the construction of the discrete gradient operator (11), the proof of Lemma 3.7 is much more technical.
  • Step $3$. To erase the term $\|w_{h}\|_{W^{1,p}(\Omega)}$ on the right hand side of (38), we prove by contradiction with the help of Lemma 3.8 which provides the discrete compactness. Then we obtain Theorem 3.2 which the stability of the FEM (14).
  • Step $4$. Roughly speaking, we apply the discrete compactness provided by Lemma 3.8 with Theorem 3.2 to prove the limit of numerical solution of the FEM (14) is a solution of (2). Then by the global $W^{2,p}$ estimate (22), the well-posedness of strong solution in $W^{2,p}(\Omega)$ of (2) is proven for the range of $p$ described in (1b). Finally by some density argument, we extend $p$ to the range described in (1a). Optimal convergence of the numerical solution of the FEM (14) is straightforward.
Theorem 3.5 and Theorem 3.6 provides well-posedness of strong solution in $W^{2,p}(\Omega)$ of (2) and optimal convergence with respect to $W_{h}^{2,p}$-norm of the numerical solution of the FEM (14), in the scenario where the coefficient matrix $A$ is discontinuous but $\gamma A$ is dominated by $I_{d}$, where the positive weight function $\gamma$ is introduced in (12). Since they are special examples of Theorem 4.1 and Theorem 4.3 respectively, the proofs of Theorem 3.5 and Theorem 3.6 are skipped. Please refer to those of Theorem 4.1 and Theorem 4.3.

### v2

@@ Abstract.
We propose one finite element method for both second order linear uniformly elliptic PDE in non-divergence form and the uniformly elliptic Hamilton-Jacobi-Bellman (HJB) equation. For both linear elliptic PDE in non-divergence form and the HJB equation, We prove the well-posedness of strong solution in $W^{2,p}(\Omega)$ and optimal convergence in discrete $W^{2,p}$-norm of the finite element approximation to the strong solution for $1<p\leq 2$ on convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$). If the domain is a two dimensional non-convex polygon, $p$ is valid in a more restricted region. Furthermore, we relax the assumptions on the continuity of coefficients of the HJB equation, which have been widely used in literature.
@@ 1. Introduction
We consider the well-posedness of the strong solution in $W^{2,p}(\Omega)$ and optimal convergence to the strong solution in discrete $W^{2,p}$-norm of a finite element method for both second order linear uniformly elliptic PDE in non-divergence form and the uniformly elliptic Hamilton-Jacobi-Bellman equation on Lipschitz polyhedral domain $\Omega$ in $\mathbb{R}^{d}$ ($d=2,3$). The Lipschitz polyhedral domain $\Omega$ is assumed to be convex if $d=3$, and can be non-convex if $d=2$.
The linear uniformly elliptic PDE in non-divergence form is find the solution $u$ satisfying
We assume that $A\in[L^{\infty}(\Omega)]^{d\times d}$ is a symmetric and uniformly positive definite matrix-valued function on $\Omega$, $\boldsymbol{b}\in[L^{\infty}(\Omega)]^{d}$, and $c\in L^{\infty}(\Omega)$ with $c\leq 0$.
The uniformly elliptic Hamilton-Jacobi-Bellman (HJB) equation is to find the solution $u$ satisfying
1.1. Literature review
There are several papers [1, 3, 6, 8, 12, 15, 19, 27, 29, 33, 32, 37] focusing on convergence to the viscosity solutions. In the Barles–Souganidis framework [1], monotonicity serves as a fundamental principle underpinning numerical schemes to ensure convergence to viscosity solutions of fully nonlinear partial differential equations. This concept is also applicable to the monotone finite difference methods for the HJB equation [3, 19]. A monotone finite element like scheme for the HJB equation was proposed in [6]. In [32], Nochetto and Zhang introduced a two-scale method based on the integro-differential formulation of (1), where a discrete Alexandroff–Bakelman–Pucci estimate was established. This two-scale method [32] was extended to the Isaacs equation [37]. However, up to our knowledge, all these numerical methods for convergence to the viscosity are of low order accuracy. In addition, most of them have stencils which increase under mesh refinement, such that the corresponding linear system may not be sparse.
With respect to numerical methods for the convergence to the strong solutions, there are three main categories based on their techniques for the linear elliptic PDEs in non-divergence form.
  • (I)
    AA can be discontinuous but $\gamma A$ dominated by $I_{d}$ where $\gamma$ is a positive weight function. In [39], Smears and Süli introduced and analyzed a Discontinuous Galerkin (DG) method for (1). The main assumptions are the coefficients $A$ is dominated by $I_{d}$ after being multiplied by a positive weight function $\gamma$ (Cordès condition) and the domain $\Omega$ is convex. The stability and optimal rate of convergence in discrete $H^{2}$-norm were proven in [39]. Following this direction, there are several papers [2, 21, 43]. The technique in [39] has been extended to the HJB equation (2) in [40, 22, 43]. It is not clear how to prove optimal rate of convergence in discrete $W^{2,p}$-norm ($p\neq 2$) for this kind of numerical methods.
  • (II)
    AA is globally continuous. In [14], Feng and Hennings and Neilan introduced and analyzed a $C^{0}$-conforming DG method for (1). The main assumptions are $A\in C^{0}(\overline{\Omega})$ and the domain $\Omega$ has $C^{1,1}$ boundary. They followed the approach in [24, Chapter $9$] to prove that their numerical method could achieve stability and optimal rate of convergence in discrete $W^{2,p}$-norm for any $1<p<+\infty$. However, this numerical method and its variant - a DG method in [17] can not deal with discontinuous coefficients $A$, since they have the jump of $A\nabla u_{h}$ along mesh interfaces ($u_{h}$ is the numerical solution). Therefore, the numerical methods in [14, 17] can not be applied for the HJB equation (2).
  • (III)
    Residual minimization methods. The main advantage of residual minimization methods [35, 41, 23] is that they can easily inherit the a priori estimate of PDE operators of (1) and (2). However, the minimal requirements of the regularity of true solution $u$ for these methods are highly affected by the choice of the norm to measure the residual in their designs. It is not clear which norm to choose for the two cases mentioned above.
A natural question is whether it is possible to design one numerical method for both linear elliptic PDE in non-divergence form and the elliptic HJB equation on Lipschitz polyhedra in $\mathbb{R}^{d}$ ($d=2,3$) and achieve optimal convergence in discrete $W^{2,p}$-norm to the strong solution with a wide range of $p$. We give positive answer in this paper.
1.2. Main contributions
With respect to the $W^{2,p}$-norm for PDE analysis and the discrete $W^{2,p}$-norm (the $W_{h}^{2,p}$-norm introduced in (7)) for numerical analysis, the ranges of $p$ considered in this paper depend on whether the the Lipschitz polyhedral domain $\Omega$ is convex. The ranges of $p$ are described below:
@@ 1.2. Main contributions
For the elliptic HJB equation (2), we further assume that there is a subset $E_{0}\subset\Omega$ with zero $d$-dimensional Lebesgue measure such that
@@
1.2.1. Contributions for the linear elliptic PDE in non-divergence form (1)
@@
When the coefficient matrix $A\in C^{0}(\overline{\Omega})$, Theorem 3.4 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ and optimal convergence with respect to $W_{h}^{2,p}$-norm (see (7)) of the numerical solution of the FEM (14). The well-posedness of (1) provided by [24, Theorem $9.15$] and optimal convergence of the FEM in [14] require $\partial\Omega$ to be $C^{1,1}$. In contrast, Theorem 3.4 is valid on Lipschitz polyhedral domains (see (5) for the detailed description). In fact, with the help of the well-posedness of strong solution stated in Theorem 3.4, the $W_{h}^{2,p}$-norm stability of the FEM in [14] can be immediately extended from domains with $C^{1,1}$ boundary to Lipschitz polyhedral domains with $p$ described in (5b). Furthermore, in contrast to the proof of numerical stability presented in [14], which rests upon a specialized duality argument (see the proof of [14, Lemma $3.5$]) contingent on the well-posedness of (1), the present work establishes numerical stability (Theorem 3.3) by invoking the discrete compactness property (Lemma 3.8) and the uniqueness of the strong solution of (1). Thus, the analysis of this paper is quite different from that of [14].
We explain the methodology to prove Theorem 3.4 by the following steps.
  • Step $1$. We prove Theorem 3.2 which give the following global $W^{2,p}$ estimate (26) of the linear elliptic PDE in non-divergence form (1):
    $\|w\|_{W^{2,p}(\Omega)}\leq C\|A:D^{2}w+\boldsymbol{b}\cdot\nabla w+cw\|_{L^{p}(\Omega)},\quad\forall w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega),$
    where the range of $p$ is described in (5a). The most difficult part of the proof of Theorem 3.2 is to obtain the above estimate for $1<p\leq\frac{3}{2}$ if $\Omega$ is a convex polyhedra in $\mathbb{R}^{3}$. It is because the Alexandroff–Bakelman–Pucci estimate can not be applied directly due to the fact $W^{2,p}(\Omega)$ can not be embedded into $C^{0}(\overline{\Omega})$. To overcome this difficulty, we provide Lemma 3.1 which shows the uniqueness of the strong solution of (1) in $W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$. Then Theorem 3.2 basically follows [34, Theorem $1.1$] by a proof by contraction.
  • Step $2$. We provide Lemma 3.7 which give the inequality (42):
    $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq C\left(\|\mathcal{L}_{\tilde{A},h}w_{h}\|_{L_{h}^{p}(\Omega)}+\|w_{h}\|_{W^{1,p}(\Omega)}\right),\qquad\forall w_{h}\in V_{h}.$
    Here, $V_{h}$ is the finite element space introduced in Section 2.1, $\mathcal{L}_{\tilde{A},h}$ is the Riesz representation (see (40)) of the FEM (14) with $\boldsymbol{b}=\boldsymbol{0}$ and $c=0$. The proof of Lemma 3.7 mimics that of [14, Lemma $3.4$]. However, due to the construction of the discrete gradient operator (11), the proof of Lemma 3.7 is much more technical.
  • Step $3$. To erase the term $\|w_{h}\|_{W^{1,p}(\Omega)}$ on the right hand side of (42), we prove by contradiction with the help of the discrete compactness - Lemma 3.8 and the uniqueness of strong solution of (1) by Theorem 3.2. Then we obtain Theorem 3.2 which the stability of the FEM (14):
    $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq C\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}\|_{L_{h}^{p}(\Omega)},\qquad\forall w_{h}\in V_{h}.$
    We would like to point out that to obtain the same result, the authors of [14] used a specialized duality argument (see the proof of [14, Lemma $3.5$]), which relies on the well-posedness of strong solution of (1).
  • Step $4$. Roughly speaking, we apply the discrete compactness - Lemma 3.8 with Theorem 3.2 to prove the limit of numerical solution of the FEM (14) is a solution of (1). Then by the global $W^{2,p}$ estimate (26), the well-posedness of strong solution in $W^{2,p}(\Omega)$ of (1) is proven for the range of $p$ described in (5b). Finally by some density argument, we extend $p$ to the range described in (5a). Optimal convergence of the numerical solution of the FEM (14) is straightforward.
@@
When the coefficient matrix $A$ is discontinuous but $\gamma A$ is dominated by $I_{d}$ (the positive weight function $\gamma$ is introduced in (12)), Theorem 3.5 and Theorem 3.6 provide well-posedness of strong solution in $W^{2,p}(\Omega)$ of (1) and optimal convergence with respect to $W_{h}^{2,p}$-norm of the numerical solution of the FEM (14). We would like to point out that Theorem 3.5 and Theorem 3.6 are special cases of Theorem 4.1 and Theorem 4.3 for the elliptic HJB equation, respectively.
1.2.2. Contributions for the elliptic HJB equation (2)
@@
For the elliptic HJB equation (2), the main theoretical results are Theorem 4.1 and Theorem 4.3. Theorem 4.1 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ of the elliptic HJB equation (2), if the coefficients satisfy the condition (83). In parallel, Theorem 4.3 provides optimal convergence of the FEM (16) in $W_{h}^{2,p}$-norm, if the coefficients satisfy the condition (93). When the domain is convex and $p=2$, the condition (83) is identical to the Cordès condition used in [40] but the condition (93) is more restrictive. Please see Remark 4.2 and Remark 4.4 for the special case $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$ and $c^{\alpha}=0$ for any $\alpha\in\Lambda$.
@@
We would like to emphasize that [40] assumes $A^{\alpha}\in C^{0}(\overline{\Omega}\times\Lambda)$ and the index set $\Lambda$ is a compact metric space, while [23, Proposition $4.3$] assumes that there are $\beta\in(0,1]$ and $0<C<+\infty$ such that $\|A^{\alpha}\|_{C^{0,\beta}(\overline{\Omega})}<C$ for any $\alpha\in\Lambda$ ([23, Proposition $4.3$] assumes $\boldsymbol{b}^{\alpha}=\boldsymbol{0}$, $c^{\alpha}=0$, and $f^{\alpha}=f$). So, if the index set $\Lambda$ has a single element, the elliptic HJB equation (2) will not become the linear elliptic PDE in non-divergence form (1) with discontinuous matrix coefficient $A$. In contrast, the assumption (6) allows discontinuity of $A^{\alpha}$, $\boldsymbol{b}^{\alpha}$ and $c^{\alpha}$. By Lemma 4.3, we can conclude that the assumptions used by [40] and [23, Proposition $4.3$] imply that the assumption (6) holds.
@@ 3. Linear elliptic PDE in non-divergence form
Theorem 3.4 provides the well-posedness of strong solution in $W^{2,p}(\Omega)$ and optimal convergence with respect to $W_{h}^{2,p}$-norm (see (7)) of the numerical solution of the FEM (14), when the coefficient matrix $A\in C^{0}(\overline{\Omega})$.
Theorem 3.5 and Theorem 3.6 provides well-posedness of strong solution in $W^{2,p}(\Omega)$ of (1) and optimal convergence with respect to $W_{h}^{2,p}$-norm of the numerical solution of the FEM (14), in the setting where the coefficient matrix $A$ is discontinuous but $\gamma A$ is dominated by $I_{d}$, where the positive weight function $\gamma$ is introduced in (12). Since they are special examples of Theorem 4.1 and Theorem 4.3 respectively, the proofs of Theorem 3.5 and Theorem 3.6 are skipped. Please refer to those of Theorem 4.1 and Theorem 4.3.